# First Light and Reionization Epoch Simulations (FLARES) - XV: The physical properties of super-massive black holes and their impact on galaxies in the early Universe

### v2

@@ Abstract
Understanding the co-evolution of super-massive black holes (SMBHs) and their host galaxies remains a key challenge of extragalactic astrophysics, particularly the earliest stages at high-redshift. However, studying SMBHs at high-redshift with cosmological simulations, is challenging due to the large volumes and high-resolution required. Through its innovative simulation strategy, the First Light And Reionisation Epoch Simulations (FLARES) suite of cosmological hydrodynamical zoom simulations allows us to simulate a much wider range of environments which contain SMBHs with masses extending to $M_{\bullet}>10^{9}\ {\rm M_{\odot}}$ at $z=5$. In this paper, we use FLARES to study the physical properties of SMBHs and their hosts in the early Universe ($5\leq\,z\leq 10$). FLARES predicts a sharply declining density with increasing redshift, decreasing by a factor of 100 over the range $z=5\to 10$. Comparison between our predicted bolometric luminosity function and pre-JWST observations yield a good match. However, recent JWST observations appear to suggest a larger contribution of SMBHs than previously observed, or predicted by FLARES. Finally, by using a re-simulation with AGN feedback disabled, we explore the impact of AGN feedback on their host galaxies. This reveals that AGN feedback results in a reduction of star formation activity, even at $z>5$, but only in the most massive galaxies. A deeper analysis reveals that AGN are also the cause of suppressed star formation in passive galaxies but that the presence of an AGN doesn’t necessarily result in the suppression of star formation.
@@ 1 Introduction
Observations of high redshift quasars (e.g. Jiang et al., 2016; Matsuoka et al., 2016; Maiolino et al., 2023b) have revealed that SMBHs existed in the Universe less than a billion years after the Big Bang (see Inayoshi et al., 2020; Fan et al., 2023, for recent reviews). The traditional stellar remnant BH formation scheme cannot explain such massive BHs in the early Universe; in order to grow BHs up to $\sim 10^{9}\ {\rm M_{\odot}}$ at $z\gtrsim 5$, alternative formation mechanisms are required, such as massive seeds and/or enhanced BH accretion (Latif et al., 2013; Volonteri et al., 2021). Suggested massive seed formation scenarios include direct collapse black holes, remnants of Population III stars, and the collapse of very massive stars formed through mergers (Loeb & Rasio, 1994; Madau & Rees, 2001; Bromm & Loeb, 2003; Portegies Zwart et al., 2004; Volonteri & Rees, 2005; Begelman et al., 2006; Regan & Haehnelt, 2009). Understanding not only the formation but also the evolution and physics of high redshift SMBHs is essential in order to capture the effects that SMBHs have on their surroundings. Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy into it (e.g. Fabian, 2012), they affect their surroundings, thus altering the overall properties of their host galaxies (e.g. the bright end of the galaxy luminosity function Kauffmann & Haehnelt, 2000; Granato et al., 2004; Bower et al., 2006; Cattaneo et al., 2006; Croton et al., 2006).
@@
From a theoretical/computational point of view, black hole physics has been an integral component of models of galaxy formation, which try to capture the effects of black hole feedback on the simulated galaxies (see the reviews of Somerville & Davé, 2015; Naab & Ostriker, 2017; Vogelsberger et al., 2020; Habouzit et al., 2022a, b). Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy to it (e.g. Fabian, 2012), they affect their surroundings thus altering the overall properties of their host galaxies. Traditionally, black hole feedback has been incorporated either through a thermal / quasar mode, where a fraction of the bolometric luminosity is injected as thermal energy to the surrounding environment (Springel et al., 2005; Booth & Schaye, 2009; Tremmel et al., 2017) or as kinetic (also known as mechanical or radio mode) mode (Croton et al., 2006; Costa et al., 2014; Choi et al., 2015; Costa et al., 2020), or as a combination of different modes (Sijacki et al., 2007; Dubois et al., 2012; Sijacki et al., 2015; Weinberger et al., 2017; Davé et al., 2019). However, different implementations of black hole physics result in discrepancies in the predictions of black hole properties both at low and at high redshifts (Meece et al., 2017; Habouzit et al., 2022a, b), which makes understanding the co–evolution of black holes and galaxies even more challenging.
@@ 2.1 FLARES
Figure 1: The number of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ (thin line) and $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ (thick line) super–massive black holes in FLARES (solid–line) and the EAGLE (dashed line) reference volume at $z=5-10$.
FLARES is a suite of 40 hydrodynamical re-simulations of spherical regions of size $\rm 14\,cMpc\,$$h^{-1}$ utilising the AGNdT9 variant of the EAGLE (Schaye et al., 2015; Crain et al., 2015) physics model and a Planck year 1 cosmology ($\rm\Omega_{M}=0.307,\;\Omega_{\Lambda}=0.693,\;$h$\;=0.6777$, Planck Collaboration et al., 2014). The regions were selected from a large (3.2 cGpc)${}^{3}$ low-resolution dark matter only (DMO) simulation (Barnes et al., 2017). The selected regions encompass a wide range of overdensities, $\delta_{14}=-0.4\to 1.0$, with greater representation at the extremes, particularly extreme over-densities. This enables us to simulate many more massive galaxies than possible using a periodic simulation and the same computational resources. For example, FLARES contains approximately 100 times as many $M_{\star}>10^{9}\ {\rm M_{\odot}}$ galaxies at $z=10$ than the fiducial EAGLE reference simulation. This makes FLARES ideally suited to studying SMBHs, and particularly AGN, since they are rare and preferentially occur in massive galaxies. This is demonstrated in Figure 1, where we show the total number of SMBHs with $M_{\bullet}/{\rm M_{\odot}}>\{10^{7},10^{8}\}$ as a function of redshift for both the (100 Mpc)${}^{3}$ EAGLE reference simulation and all FLARES regions combined. At $z=5$ FLARES contains 6 (20) times as many SMBHs with $M_{\bullet}>10^{7}(10^{8})\ {\rm M_{\odot}}$, respectively. This disparity increases with increasing redshift: there are no SMBHs with $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ in EAGLE at $z>6$, for example, whereas those in FLARES extend to $z=9$.
@@ 2.1.1 Weighting
An important consequence of the FLARES strategy is that universal cosmological scaling relations and distribution functions (e.g. the SMBH mass function) cannot be trivially recovered. Instead, it is necessary to weight each simulation/region by how likely it is to occur in the parent DMO simulation. This is described in more detail in Lovell et al. (2021) where we demonstrate its appropriateness by recovering the galaxy stellar mass function. In brief, we measure the overdensity in the parent box on a $2.67\;\mathrm{Mpc}$ grid, find the distribution of overdensities within each region, and then weight each region $j$ by some factor
$f_{j}=\sum_{i}r_{i}w_{ij}\;\;,$ (1)
where $r_{i}$ is the ratio of the true weighting in overdensity bin $i$ (over the whole parent box) to the summed weight of all regions in that overdensity bin, and $w_{ij}$ is the weighting of region $j$ in overdensity bin $i$.
@@ 2.2 Black Hole Modelling in EAGLE
In this section we summarise the SMBH physics utilised by the EAGLE model which is employed by FLARES. For a full description of FLARES see Lovell et al. (2021); Vijayan et al. (2021), and for the EAGLE model see Schaye et al. (2015) and Crain et al. (2015). In short, BHs in the EAGLE model are seeded into sufficiently massive halos and then allowed to grow through accretion and mergers. A fraction of the rest-mass accreted onto the disc is able to be radiated away. A fraction of radiated energy is injected into neighbouring gas particles, heating them.
@@ 2.2.1 Seeding
BHs are seeded into halos exceeding a halo mass of $\log_{10}(M_{\rm h}/h^{-1}{\rm M_{\odot}})=10$ by converting the highest density gas particle into a BH particle (Springel et al., 2005). BHs carry both a particle mass and a subgrid mass. The particle initial mass is set by that of the converted gas particle while the initial subgrid mass is $\log_{10}(M_{\bullet}/h^{-1}{\rm M_{\odot}})=5$. The use of a separate subgrid mass is necessary because the black hole seed mass is below the simulation mass resolution. Calculations pertaining to growth and feedback events of the black hole are computed using the subgrid mass, $M_{\bullet}$, while gravitational interactions use the particle mass. When the sub-grid BH mass exceeds the particle mass the BH particle is allowed to stochastically accrete a neighbouring gas particle. When the sub-grid BH mass is much larger than the gas particle mass, the BH sub-grid and particle masses effectively grow together. The seed and gas particle masses place an effective lower-limit on the BH masses which are robust, or resolved. For this analysis we conservatively assume that BHs are considered resolved at $M_{\bullet}=10^{7}\ {\rm M_{\odot}}$ and focus our analysis on these objects.
@@ 2.2.2 Accretion
The primary growth of BHs is through accretion. The EAGLE subgrid accretion model allows the BH to accrete material at a maximum rate determined by the Eddington accretion rate scaled by a factor of $1/h$ (McAlpine et al., 2020), i.e.
@@
where $G$ is the gravitational constant, $m_{\bullet}$ is the black hole subgrid mass, $m_{\rm p}$ is the proton mass, $\epsilon_{\rm r}$ is the radiative efficiency of the accretion disk, $\sigma_{\rm T}$ is the Thomson cross-section and $c$ is the speed of light. The radiative efficiency is set to $\epsilon_{\rm r}=0.1$.
The floor for the accretion rate is set by the Bondi–Hoyle accretion rate defined as
@@
where $c_{\rm s}$ and $v$ are the speed of sound and the relative velocity of the BH and gas, respectively (Bondi & Hoyle, 1944). Hence, the actual accretion rate onto the black hole is then given by
where $C_{\rm visc}$ is a parameter related to the viscosity of the accretion disc (see next subsection) and $V_{\upphi}$ is the rotation speed of the gas around the BH (see equation 16 of Rosas-Guevara et al., 2015). To re-iterate, the maximum of the accretion rate is set by equation 2. Finally, this results in the BH mass growth rate of
@@
The accretion rates of each BH are reported at every time–step the BH is active and, for all BHs, at the snapshot redshift. The rates reported in a single time–step can differ significantly from rates averaged across longer timescales. In Appendix A we contrast the instantaneous accretion rates with those averaged on longer timescales.
@@ 2.2.3 Emission
The total (bolometric) luminosity radiated by an AGN is simply proportional to the accretion disc or BH growth rate as,
@@
In this work we only present this bolometric luminosity. However, in a companion work (Wilkins et al. in-prep) we fully model the spectral energy distributions, including the contribution of nebular continuum and line emission, of AGN using a subgrid model combining the disc, narrow and broad line regions, and a dusty torus.
@@ 2.2.4 Mergers
In addition to accretion, BHs can also grow by merging. BHs are merged if their separation is smaller than three gravitational softening lengths and are within the BH smoothing kernel length, $h_{\bullet}$, of each other and have a relative velocity of
@@
where $m_{\bullet}$ is the subgrid mass of the bigger of the merging BHs. The limiting velocity given by equation 7 prevents BH mergers during the initial stages of galaxy mergers and makes them only possible once some time of the initial merger has passed and the relative velocities have settled. As a consequence galaxies in the EAGLE model can often host multiple black holes. We explore this in the context of FLARES in §3.1.
2.2.5 Repositioning
@@ 2.2.6 Feedback
In terms of the subgrid AGN feedback model FLARES adopts the modelling approach of the C-EAGLE simulations (Barnes et al., 2017), using the AGNdT9 subgrid parameter configuration. This configuration is parameterised with $C_{\mathrm{visc}}=2\uppi\times 10^{2}$ and $\Delta T_{\rm AGN}=10^{9}\,$, where the former is a free parameter controlling the sensitivity of the BH accretion rate to the angular momentum of the gas and the latter is the temperature increase of gas during AGN feedback.
AGN feedback in the EAGLE model is modelled with only one feedback channel, contrasting with the multi-mode feedback implemented in other simulations including Simba (Davé et al., 2019) and TNG (Zinger et al., 2020). In this model, thermal energy is injected stochastically to gas particles neighbouring the BH particle, in a kernel–weighted manner. The energy injection rate is calculated as a fraction of the total accretion rate, as $\epsilon_{\rm f}\epsilon_{\rm r}\dot{m}_{\rm accr}c^{2}$, where $\epsilon_{\rm f}=0.15$ is the fraction of the feedback energy that is coupled to the ISM and $\epsilon_{\rm r}=0.1$ is the radiative efficiency of the accretion processes introduced in Eq. 5.
The probability of injecting energy into each nearby gas particle is given by
@@
where $E_{\bullet}$ represents a ‘reservoir’ of feedback energy carried by the black hole, with $\epsilon_{\rm r}\epsilon_{\rm f}\dot{m}_{\rm accr}c^{2}\Delta t$ added to it after every time step $\Delta t$, $\Delta\epsilon_{\rm AGN}$ is the change in internal energy per unit mass of a gas particle corresponding to a temperature increase of $\Delta T_{\rm AGN}$, $N_{\mathrm{ngb}}$ is the number of gas neighbours of the BH and $\left\langle m_{\rm g}\right\rangle$ is their mean mass. The reservoir, $E_{\bullet}$, is then decreased by the amount of the injected energy.
Since $\Delta T_{\rm AGN}$ directly determines the amount of energy in each AGN feedback event, it is the most important factor in modelling the feedback from BH accretion. A larger value results in more energy dumped into the neighbouring particles of the BH, but also makes the feedback events more rare, as the change in internal energy of a gas particle is directly proportional to temperature increase and the probability of energy injection into a gas particle is inversely proportional to the internal energy increase. The temperature increase from AGN feedback was set to a value higher than that from stellar feedback due to gas densities being higher in the vicinity of black holes than they are for typical star-forming gas (Crain et al., 2015).

### v3

@@ Abstract
Understanding the co-evolution of super-massive black holes (SMBHs) and their host galaxies remains a key challenge of extragalactic astrophysics, particularly the earliest stages at high-redshift. However, studying SMBHs at high-redshift with cosmological simulations, is challenging due to the large volumes and high-resolution required. Through its innovative simulation strategy, the First Light And Reionisation Epoch Simulations (FLARES) suite of cosmological hydrodynamical zoom simulations allows us to simulate a much wider range of environments which contain SMBHs with masses extending to $M_{\bullet}>10^{9}\ {\rm M_{\odot}}$ at $z=5$. In this paper, we use FLARES to study the physical properties of SMBHs and their hosts in the early Universe ($5\leq\,z\leq 10$). FLARES predicts a sharply declining density with increasing redshift, decreasing by a factor of 100 over the range $z=5\to 10$. Comparison between our predicted mass and bolometric luminosity functions and pre-JWST observations yielded a reasonable match. However, recent JWST observations appear to suggest a higher density of SMBHs at $z\approx 5$ and the presence of more luminous SMBHs at $z>6$ than predicted by FLARES. Finally, by using a re-simulation with AGN feedback disabled, we explore the impact of AGN feedback on their host galaxies. This reveals that AGN feedback results in a reduction of star formation activity, even at $z>5$, but only in the most massive galaxies. A deeper analysis reveals that AGN are also the cause of suppressed star formation in passive galaxies but that the presence of an AGN doesn’t necessarily result in the suppression of star formation.
@@ 1 Introduction
Observations of high redshift quasars (e.g. Jiang et al., 2016; Matsuoka et al., 2016; Maiolino et al., 2023b) have revealed that SMBHs existed in the Universe less than a billion years after the Big Bang (see Inayoshi et al., 2020; Fan et al., 2023, for recent reviews). The traditional Eddington-limited stellar remnant BH formation scheme cannot explain such massive BHs in the early Universe; in order to grow BHs up to $\sim 10^{9}\ {\rm M_{\odot}}$ at $z\gtrsim 5$, alternative formation mechanisms are required, such as massive seeds and/or enhanced BH accretion (Latif et al., 2013; Volonteri et al., 2021). Suggested massive seed formation scenarios include direct collapse black holes and the collapse of very massive stars formed through mergers (Loeb & Rasio, 1994; Madau & Rees, 2001; Bromm & Loeb, 2003; Portegies Zwart et al., 2004; Volonteri & Rees, 2005; Begelman et al., 2006; Regan & Haehnelt, 2009). Understanding not only the formation but also the evolution and physics of high redshift SMBHs is essential in order to capture the effects that SMBHs have on their surroundings. Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy into it (e.g. Fabian, 2012), they affect their surroundings, thus altering the overall properties of their host galaxies (e.g. the bright end of the galaxy luminosity function Kauffmann & Haehnelt, 2000; Granato et al., 2004; Bower et al., 2006; Cattaneo et al., 2006; Croton et al., 2006).
@@
From a theoretical/computational point of view, black hole physics has been an integral component of models of galaxy formation, which try to capture the effects of black hole feedback on the simulated galaxies (see the reviews of Somerville & Davé, 2015; Naab & Ostriker, 2017; Vogelsberger et al., 2020; Habouzit et al., 2022a, b). Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy to it (e.g. Fabian, 2012), they affect their surroundings thus altering the overall properties of their host galaxies (Ciotti & Ostriker, 2001; Di Matteo et al., 2005; Murray et al., 2005; Hopkins et al., 2006). Traditionally, black hole feedback has been incorporated either through a thermal or quasar mode, where a fraction of the bolometric luminosity is injected as thermal energy to the surrounding environment (Springel et al., 2005; Booth & Schaye, 2009; Tremmel et al., 2017) or as kinetic mode (Croton et al., 2006; Costa et al., 2014; Choi et al., 2015; Costa et al., 2020), or as a combination of different modes (Sijacki et al., 2007; Dubois et al., 2012; Sijacki et al., 2015; Weinberger et al., 2017; Davé et al., 2019). However, different implementations of black hole physics result in discrepancies in the predictions of black hole properties both at low and at high redshifts (Meece et al., 2017; Habouzit et al., 2022a, b), which makes understanding the co–evolution of black holes and galaxies even more challenging.
@@ 2.1 FLARES
Figure 1: The number of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ (thin line) and $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ (thick line) super–massive black holes in FLARES (solid–line) and the EAGLE (dashed line) reference volume at $z=5-10$. The FLARES simulation strategy results in the simulating of many more massive SMBHs than the original EAGLE reference simulation despite a comparable total volume.
The core FLARES simulations are a suite of 40 independent hydrodynamical re-simulations of spherical regions of size $\rm 14\,cMpc\,$$h^{-1}$ utilising the GADGET–3 code (see e.g. Springel, 2005; Springel et al., 2021), the AGNdT9 variant of the EAGLE (Schaye et al., 2015; Crain et al., 2015) physics model, and a Planck year 1 cosmology ($\rm\Omega_{M}=0.307,\;\Omega_{\Lambda}=0.693,\;$h$\;=0.6777$, Planck Collaboration et al., 2014). FLARES adopts an identical resolution to the fiducial EAGLE simulation, with dark-matter (DM) and gas particle masses of $m_{dm}=9.7\times 10^{6}\ {\rm M_{\odot}}$ and $m_{g}=1.8\times 10^{6}\ {\rm M_{\odot}}$ respectively, and a softening length of $2.66\ {\rm ckpc}$. The EAGLE model uses a list of 11 elements from Wiersma et al. (2009) to calculate the radiative cooling and photoheating rates. Hydrogen reionisation is implemented as a uniform, time–dependent ionizing background (Haardt & Madau, 2001) that begins at $z$ = 11.5. Finally, star formation is modeled stochastically following the prescription of Dalla Vecchia & Schaye (2008); Schaye (2004), coupled with a metallicity-dependent density threshold; and the resulting stellar particles are subject to mass loss and type Ia supernovae.
The 40 FLARES regions were selected from a large (3.2 cGpc)3 low-resolution dark matter only (DMO) simulation (Barnes et al., 2017). The selected regions encompass a wide range of overdensities, $\delta_{14}=-0.4\to 1.0$, with greater representation at the extremes, particularly extreme over-densities. This enables us to simulate many more massive galaxies than possible using a periodic simulation and the same computational resources. For example, FLARES contains approximately 100 times as many $M_{\star}>10^{9}\ {\rm M_{\odot}}$ galaxies at $z=10$ than the fiducial EAGLE reference simulation, despite simulating a comparable total volume. This makes FLARES ideally suited to studying SMBHs, and particularly AGN, since they are rare and preferentially occur in massive galaxies. This is demonstrated in Figure 1, where we show the total number of SMBHs with $M_{\bullet}/{\rm M_{\odot}}>\{10^{7},10^{8}\}$ as a function of redshift for both the (100 Mpc)3 EAGLE reference simulation and all FLARES regions combined. At $z=5$ FLARES contains 6 (20) times as many SMBHs with $M_{\bullet}>10^{7}(10^{8})\ {\rm M_{\odot}}$, respectively. This differences increases with increasing redshift: there are no SMBHs with $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ in EAGLE at $z>6$, for example, whereas those in FLARES extend to $z=9$. It is important to stress however that while FLARES simulates more massive SMBHs than the EAGLE reference simulation the predicted number densities are consistent once the individual simulations are appropriately weighted (see §2.1.1, and §3.2). In addition to the core suite several regions have been run with model variations. For example, for several regions we “turned off” AGN feedback to explore its impact on galaxy properties; this is discussed in §4.4.
@@ 2.1.1 Weighting
An important consequence of the FLARES strategy is that universal cosmological scaling relations and distribution functions (e.g. the SMBH mass function) cannot be trivially recovered. Instead, it is necessary to weight each simulation/region by how likely it is to occur in the parent DMO simulation. This is described in more detail in Lovell et al. (2021) where we demonstrate its appropriateness by recovering the galaxy stellar mass function.
In short, the $(3.2\ \rm{cGpc})^{3}$ parent volume is split into grid cells, $2.67\ \rm{cMpc}$ in length. The overdensity in each grid cell is recorded, and the distribution of overdensities is divided into 50 equal-width bins. Each re-simulated region, $r$, is assigned a weight, $w_{r}$, given by
$w_{r}=\sum_{i}f_{r,i}\frac{f_{\rm{parent},\mathit{i}}}{f_{\rm{resims},\mathit{i}}},$ (1)
where $f_{r,i}$ is the fraction of grid cells in region $r$ that occupy overdensity bin $i$, $f_{\rm{parent},\mathit{i}}$ is the fraction of grid cells in the complete parent simulation that fall within overdensity bin $i$, and $f_{\rm{resims},\mathit{i}}$ is the fraction of grid cells in all 40 resimulated regions that are contained within overdensity bin $i$. Thus, the weight carried by a region is proportional to how representative that region is of the universal mean.
@@ 2.2 Black Hole Modelling in EAGLE
In this section we summarise the SMBH physics utilised by the EAGLE model which is employed by FLARES. For a full description of FLARES see Lovell et al. (2021); Vijayan et al. (2021), and for the EAGLE model see Schaye et al. (2015) and Crain et al. (2015). In short, BHs in the EAGLE model are seeded into sufficiently massive halos and then allowed to grow through accretion and mergers. A fraction of the rest-mass accreted onto the SMBH is able to be radiated away. A fraction of radiated energy is injected into neighbouring gas particles, heating them.
@@ 2.2.1 Seeding
BHs are seeded into halos exceeding a halo mass of $M_{\rm h}=10^{10}\ h^{-1}{\rm M_{\odot}}$ by converting the highest density gas particle into a SMBH particle (Springel et al., 2005). BHs carry both a particle mass and a subgrid mass. The particle initial mass is set by that of the converted gas particle while the initial subgrid mass is $M_{\rm\bullet,\ seed}=10^{5}\ h^{-1}{\rm M_{\odot}}$. The use of a separate subgrid mass is necessary because the black hole seed mass is below the simulation mass resolution. Calculations pertaining to growth and feedback events of the black hole are computed using the subgrid mass, $M_{\bullet}$, while gravitational interactions use the particle mass. When the sub-grid SMBH mass exceeds the particle mass the SMBH particle is allowed to stochastically accrete a neighbouring gas particle. When the sub-grid SMBH mass is much larger than the gas particle mass, the SMBH sub-grid and particle masses effectively grow together. The seed and gas particle masses place an effective lower-limit on the SMBH masses which are robust, or resolved. For this analysis we conservatively assume that BHs are considered resolved at $M_{\bullet}=10^{7}\ {\rm M_{\odot}}$ and focus our analysis on these objects.
@@ 2.2.2 Accretion
The primary growth of SMBHs is through accretion. The EAGLE subgrid accretion model allows the SMBH to accrete material at a maximum rate determined by the Eddington accretion rate scaled by a factor of $1/h$ (McAlpine et al., 2020), i.e.
@@
where $G$ is the gravitational constant, $M_{\bullet}$ is the black hole subgrid mass, $m_{\rm p}$ is the proton mass, $\epsilon_{\rm r}$ is the radiative efficiency of the accretion disk, $\sigma_{\rm T}$ is the Thomson cross-section and $c$ is the speed of light. The radiative efficiency in FLARES/EAGLE is set to $\epsilon_{\rm r}=0.1$.
The floor for the accretion rate is set by,
where $C_{\rm visc}$ is a parameter related to the viscosity of the accretion disc (see next subsection) and $V_{\upphi}$ is the rotation speed of the gas around the SMBH (see equation 16 of Rosas-Guevara et al., 2015) and $\dot{M}_{\rm Bondi}$ is the Bondi–Hoyle accretion rate defined as,
@@
where $c_{\rm s}$ and $v$ are the speed of sound and the relative velocity of the SMBH and gas, respectively (Bondi & Hoyle, 1944).
Finally, this results in the SMBH mass growth rate of
@@
The accretion rates of each SMBH are reported at every time–step the SMBH is active and, for all BHs, at the snapshot redshift. Due to numerical noise the rates reported in a single time–step can differ significantly from rates averaged across longer timescales as discussed in §3.3.1.
@@ 2.2.3 Emission
The total (bolometric) luminosity radiated by an AGN is simply proportional to the accretion disc or SMBH growth rate as,
@@ 2.2.4 Mergers
In addition to accretion, SMBHs can also grow by merging. SMBHs are merged if their separation is smaller than three gravitational softening lengths and are within the SMBH smoothing kernel length, $h_{\bullet}$, of each other and have a relative velocity of
@@
where $M_{\bullet}$ is the subgrid mass of the bigger of the merging SMBHs. The limiting velocity given by equation 7 prevents SMBH mergers during the initial stages of galaxy mergers and makes them only possible once some time of the initial merger has passed and the relative velocities have settled. As a consequence galaxies in the EAGLE model can often host multiple black holes. We explore this briefly in the context of FLARES in §3.1.
2.2.5 Dynamical Fraction
@@ 2.2.6 Feedback
AGN feedback in the EAGLE model is modelled with only one feedback channel, contrasting with the multi-mode feedback implemented in other simulations including Simba (Davé et al., 2019) and TNG (Zinger et al., 2020). In this model, described fully in Schaye et al. (2015) and based on Booth & Schaye (2009), thermal energy is injected stochastically to gas particles neighbouring the SMBH particle, in a kernel–weighted manner.
The energy injection rate is calculated as a fraction of the total accretion rate, as $\epsilon_{\rm f}\epsilon_{\rm r}\dot{M}_{\rm accr}c^{2}$, where $\epsilon_{\rm f}=0.15$ is the fraction of the feedback energy that is coupled to the ISM and $\epsilon_{\rm r}=0.1$ is the radiative efficiency of the accretion processes introduced in Eq. 5. Every time-step energy ($\Delta E=\epsilon_{\rm r}\epsilon_{\rm f}\dot{M}_{\rm accr}c^{2}\Delta t$) is added to reservoir of feedback energy $E_{\bullet}$. If the SMBH has sufficient stored energy to heat at least $n_{\rm heat}$ particles (where here we assume $n_{\rm heat}=1$) of mass $m_{g}$ by $\Delta T_{\rm AGN}$ then the SMBH is allowed to stochastically heat each of its neighbouring particles by increasing their temperature by $\Delta T_{\rm AGN}$. The probability of injecting energy into each nearby gas particle is given by,
@@
where $\Delta\epsilon_{\rm AGN}$ is the change in internal energy per unit mass of a gas particle corresponding to a temperature increase of $\Delta T_{\rm AGN}$, $N_{\mathrm{ngb}}$ is the number of gas neighbours of the SMBH and $\left\langle m_{\rm g}\right\rangle$ is their mean mass. The reservoir, $E_{\bullet}$, is then decreased by the amount of the injected energy.
Since $\Delta T_{\rm AGN}$ directly determines the amount of energy in each AGN feedback event, it is the most important factor in modelling the feedback from SMBH accretion. A larger value results in more energy dumped into the neighbouring particles of the BH, but also makes the feedback events more rare, as the change in internal energy of a gas particle is directly proportional to temperature increase and the probability of energy injection into a gas particle is inversely proportional to the internal energy increase. The temperature increase from AGN feedback was set to a value higher than that from stellar feedback due to gas densities being higher in the vicinity of black holes than they are for typical star-forming gas (Crain et al., 2015). In FLARES we adopt the approach of the C-EAGLE simulations (Barnes et al., 2017), using the AGNdT9 subgrid parameter configuration. This configuration is parameterised with $C_{\mathrm{visc}}=2\uppi\times 10^{2}$ and $\Delta T_{\rm AGN}=10^{9}\,$, where the former is a free parameter controlling the sensitivity of the SMBH accretion rate to the angular momentum of the gas and the latter is the temperature increase of gas during AGN feedback.
2.3 Stellar Emission
In Section 4 we make comparisons between the bolometric luminosities of the SMBHs and the stellar content of galaxies. The spectral energy distribution modelling in FLARES is described in depth in Vijayan et al. (2021); in short, every star particle in each galaxy is associated with an SED based on its mass, age, and metallicity assuming a particular choice of stellar population synthesis model and initial mass function (IMF). In this work we assume v2.2.1 of the Binary Population and Spectral Synthesis (BPASS) model (Stanway & Eldridge, 2018) and a Chabrier (2003) IMF. While in Vijayan et al. (2021) we also consider the impact of reprocessing by dust and gas here we ignore those effects since we are only interested in the bolometric output.